40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Tue Jun 23 20:37:22 2015, Evan, DailyProgress, optic, Pointing recovered 
    Reply  Wed Jun 24 16:19:37 2015, Evan, DailyProgress, optic, Locking recovered 
Message ID: 1553     Entry time: Wed Jun 24 16:19:37 2015     In reply to: 1552
Author: Evan 
Type: DailyProgress 
Category: optic 
Subject: Locking recovered 

I reinstalled the old, underpowered unipolar HV supply that we used to use for the south cavity. Since Aidan is going to redo the power distribution anyway, there's no point in fussing with it now. The south cavity locks fine. The digital temperature offloading seems to be working as well. Light incident on each cavity is about 5 mW.

ELOG V3.1.3-